help with problem:
m-6(m+10)=50
Need all steps. Thanks.

Answer by midwood_trail(310)


m-6(m+10)=50
First thing to do is remove the parentheses using the distributive rule.

m - 6(m + 10) = 50
m - 6m -60 = 50
Combine like terms on the left side of the equation:
-5m - 60 = 50
-5m = 50 + 60
-5m = 110
m = 110/-5
m = -22
Did you follow?